OpenStudy (anonymous):

How did I do? 2nd order DE \[25y''+9y=0\] \[r(r+9)=0\] so r=0 and r=-9 \[y=c_2e^{-9x}\]

OpenStudy (zzr0ck3r):

im confused on what you are asking

OpenStudy (anonymous):

The question in my book reads: Solve the differential equation. I think I did that correctly, but let me know if I made any mistakes.

OpenStudy (anonymous):

Uhh... yeah, sorry, it should be \[ c_1e^{\frac{3}{5}ix}+c_2e^{-\frac{3}{5}ix}=y(x)\]

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Characteristic polynomial (roots of)

OpenStudy (anonymous):

duh. I wrote it wrong in my notebook....I solved r^2+9r=0 and it should have been 25r^2+9r

OpenStudy (anonymous):

I recommend \(25r^2+9=0\) ... I think you'd benefit from checking your work, again.

OpenStudy (anonymous):

yeah. sorry for my dumb little mistake :P

OpenStudy (anonymous):

It's all right, it happens.

OpenStudy (anonymous):

I still get a different answer than you did. \[25r^2+9r=0\] \[r(25r+9)=0\] so r=0 and \[r=\frac{-9}{25}\] \[y=c_2e^{\frac{-9}{25}x}\]

OpenStudy (anonymous):

@LolWolf ?

OpenStudy (anonymous):

Wait, how? We take: \[ 25r^2e^{rx}+9e^{rx}=0 \]And factor \(e^{rx}\) out, which we know cannot be zero, thus: \[ (25r^2+9)e^{rx}=0 \implies (25r^2+9)=0 \]

OpenStudy (anonymous):

(Remember you don't differentiate \(y(x)\) ...! Important)

OpenStudy (anonymous):

Is my auxiliary equation wrong? \[r(25r+9)=0\]

OpenStudy (anonymous):

Yeah, I think you're differentiating \(y(x)\) as the polynomial should be \(25r^2+9=0\)

OpenStudy (anonymous):

See question 6

OpenStudy (anonymous):

Ooooh...I see what I did. the auxiliary equation is \[r^2+9=0\]

OpenStudy (anonymous):

I can't read today LOL :P

OpenStudy (anonymous):

Oh gawd! \[25r^2+9=0\]

OpenStudy (anonymous):

Yes, here we have: \[ 25\frac{d^2y(x)}{dx^2}+9y(x)=0 \\ y(x)=e^{rx}\\ 25\frac{d^2e^{rx}}{dx^2}+9e^{rx}=0 \implies\\ 25r^2e^{rx}+9e^{rx}=0 \implies\\ (25r^2+9)e^{rx}=0 \implies\\ 25r^2+9=0 \]Et al, you know the drill.

OpenStudy (anonymous):

thank you. Thanks for patient so patient with me. That makes more sense now

OpenStudy (anonymous):

I can't write today either. I meant thanks for being soo patient with me. I worked the night shift, my brain is not quite up to speed today. :P

OpenStudy (anonymous):

Sure thing, glad to help. And, it's all right.